pooja195 (pooja195):

This is an insane question more out of curiosity but is it possible to see molecules and bacteria with the naked eye?

pooja195 (pooja195):

in some way I think you can because if u ever look at the sky on a gray day from a window u see little balls of light moving you might think its insane but is it possible?

OpenStudy (anonymous):

1. Finite size of pupil sets upper limit on eye's resolution The resolution of the eye is the smallest object the eye can see. This is limited by the diffraction limit, which is approximated by the formula, \theta\approx1.22\times\frac{\lambda}{D} , where \theta is the angular size of the object , \lambda is the wavelength of visible light and D is the diameter of the pupil. The angular size is just the ratio of the object's size versus the distance to the object. The normal pupil size of a human eye is 4mm, which sets a minimum angular resolution of the eye at 2\times10^{-4}rad. Obviously, we want to put small objects as close to our eyes as possible to be able to see them, but there is a minimum distance for comfortable viewing which is roughly at 25cm. This works out to a size of 0.04mm if the human eye is diffraction-limited, but unfortunately, our eyes do not work at this higher limit. However, a quoted figure for the smallest resolvable size is 0.1mm, just double the figure estimated here, showing that the diffraction limit is a crucial factor in visual resolving power. 2. Separation of cones at the back of the human eye This is mentioned by Daniel in one of the answer. Light must hit separate cones for our brain to interpret them as coming from two different 'dots'. The diameter of the human eye is about 25mm and the separation of human cones is 2µm. Again, we can calculate the angular resolution (8\times10^{-5}rad) and use the minimum comfortable viewing distance of 25cm to determine the smallest resolvable object, which is now estimated to be 0.02mm. Well... close enough to the quoted literature value of 0.1mm. Side note: No point to evolve higher-density cones Why didn't humans evolve to have more compact cones? This is because there is no point in doing so, as our vision is constrained by the diffraction limit of the pupil. From point 1, the angular resolution of 2\times10^{-4}rad corresponds to a separation of 5µm at the back of our eyes. Notice how this limit (determined from the front of the eyes at the pupil) is so close to the physical size of the cones at the back of our eyes. There will be no enhancement to our vision even if we had better cones of a smaller width! The light from a tiny object will be smeared across multiple cones at the back of our eyes and our brain will still see 2 closely-spaced points as one fuzzy point. If we need to see smaller objects more clearly, we will need to evolve both bigger eyeballs AND smaller and more compact cones at the same time. What is 0.1mm? It turns out that our visual resolution is sufficient for us to discern the largest of bacteria. Answering the question directly, we can see the largest bacteria known to us, the Thiomargarita namibiensis. This bacteria has a usual size of 0.1 to 0.3mm and is shown as the glowing white blobs above a fruit fly in the following picture. For reference, I have also included a macroscopic picture of a fruit fly which usually spans 3 to 4mm. Actually, the eye is not that great. This performance is known as the quantum efficiency of the eye and it is quoted to be 1% to 4%. This means that out of 100 times a single photon hits your photoreceptors, your brain will only realize the event about 4 times. To look at atoms you need atomic force microscopy.

OpenStudy (anonymous):

www.quora.com/Vision-eyesight/What-is-the-smallest-thing-a-human-eye-can-see-and-why

pooja195 (pooja195):

thanks!!